# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ Abstract
The Langevin Dynamics framework, which aims to generate samples from the score function of a probability distribution, is widely used for analyzing and interpreting score-based generative modeling. While the convergence behavior of Langevin Dynamics under unimodal distributions has been extensively studied in the literature, in practice the data distribution could consist of multiple distinct modes. In this work, we investigate Langevin Dynamics in producing samples from multimodal distributions and theoretically study its mode-seeking properties. We prove that under a variety of sub-Gaussian mixtures, Langevin Dynamics is unlikely to find all mixture components within a sub-exponential number of steps in the data dimension. To reduce the mode-seeking tendencies of Langevin Dynamics, we propose Chained Langevin Dynamics, which divides the data vector into patches of constant size and generates every patch sequentially conditioned on the previous patches. We perform a theoretical analysis of Chained Langevin Dynamics by reducing it to sampling from a constant-dimensional distribution. We present the results of several numerical experiments on synthetic and real image datasets, supporting our theoretical results on the iteration complexities of sample generation from mixture distributions using the chained and vanilla Langevin Dynamics. The code is available at https://github.com/Xiwei-Cheng/Chained_LD.
@@ 1 Introduction
Langevin dynamics, a.k.a. Langevin Monte Carlo, is a well-known Markov Chain Monte Carlo (MCMC) sampling methodology that has been widely used to implement and interpret score-based generative modeling. It can produce samples from the (Stein) score function of a probability density, i.e., the gradient of the log probability density function with respect to data. It has been widely recognized that a pitfall of Langevin dynamics is its slow mixing rate [1, 2, 3]. Specifically, [4] shows that under a multi-modal data distribution, the samples from Langevin dynamics may have an incorrect relative density across the modes. Based on this finding, [4] proposes anneal Langevin dynamics, which injects different levels of Gaussian noise into the data distribution and samples with Langevin dynamics on the perturbed distribution. While outputting the correct relative density across modes can be challenging for Langevin dynamics, a natural question is whether Langevin dynamics would be able to find all the modes of a multi-modal distribution.
In this work, we study this question by analyzing the mode-seeking properties of Langevin dynamics. The notion of mode-seekingness [5, 6, 7, 8] refers to the property that a generative model captures only a subset of the modes of a multi-modal distribution. We note that a similar problem, known as metastability, has been studied in the context of Langevin diffusion, a continuous-time version of Langevin dynamics described by stochastic differential equation (SDE) [9, 10, 11]. Specifically, [9] gave a sharp bound on the mean hitting time of Langevin diffusion and proved that it may require exponential (in the space dimensionality $d$) time for transition between modes. Regarding discrete Langevin dynamics, [3] constructed a probability distribution whose density is close to a mixture of two well-separated isotropic Gaussians, and proved that Langevin dynamics could not find one of the two modes within an exponential number of steps. However, further exploration of the mode-seeking tendencies of Langevin dynamics and its variants for general distributions is still lacking in the literature.
In this work, we study Langevin dynamics under multi-modal distributions in a slightly different setting from the standard theory literature on sampling. As illustrated in Figure 1, the existing theoretical literature commonly considers a mixture of well-separated modes with bounded variance. On the other hand, in our analysis, we consider a low-density high-variance mode (referred to as Mode 0 or $P^{(0)}$) surrounding the other modes and filling up the space between the modes. Note that Langevin dynamics relies on the score function (i.e., the gradient of log-pdf) to search for the modes. For a mixture model with no in-between modes (Figure 1.a), Langevin dynamics could carry information about the direction towards the closest mode, even if they are far from all the modes. However, assuming an in-between mode $P^{(0)}$ with high variance (Figure 1.b), the gradient information outside the support set of the low-variance modes will be dominated by $P^{(0)}$, despite the minor overall mass of $P^{(0)}$. As a result, one can expect that the dynamics would randomly explore a large volume in ${\mathbb{R}}^{d}$ until finding a low-variance mode, which can take a significant time.
To theoretically formulate and demonstrate the potential mode-seeking tendency of Langevin dynamics, we begin by analyzing the convergence for mixture distributions of Gaussian modes, under which Langevin dynamics could fail to visit all the mixture components within sub-exponential steps (in the data dimension). Subsequently, we generalize this result to mixture distributions of sub-Gaussian modes. This generalization extends our earlier result on Gaussian mixtures to a significantly larger family of mixture models, as the sub-Gaussian family includes any distribution over an $\ell_{2}$-norm-bounded support set.
(a) Traditional mixture model
(b) Mixture model with in-between mode $P^{(0)}$
Figure 1: Traditional mixture models studied in the literature vs. our analyzed mixture distribution possessing the in-between mode $P^{(0)}$. $P^{(0)}$ is supposed to contain a minor probability mass, yet with a significantly higher variance than the other modes $P^{(1)},\ldots,P^{(k)}$.
To reduce Langevin dynamics’ large iteration complexity shown under a high-dimensional input vector, we propose Chained Langevin Dynamics (Chained-LD). Since Langevin dynamics could suffer from the curse of dimensionality, we decompose the sample ${\mathbf{x}}\in{\mathbb{R}}^{d}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(d/Q)}$, each of constant size $Q$, and sequentially generate every patch ${\mathbf{x}}^{(q)}$ for all $q\in[d/Q]$ statistically conditioned on previous patches, i.e., $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(0)},\cdots{\mathbf{x}}^{(q-1)})$. The combination of all patches generated from the conditional distribution faithfully follows the probability density $P({\mathbf{x}})$ due to the chain rule, while drawing samples from each patch requires less cost due to the reduced dimension. We also provide a theoretical analysis of Chained-LD by reducing the convergence of a $d$-dimensional sample to the convergence of each patch.
@@
Finally, we present the results of several numerical experiments to validate our theoretical findings. For synthetic experiments, we consider moderately high-dimensional Gaussian mixture models, where the vanilla Langevin dynamics could not find all the components within a million steps, while Chained-LD could capture all the components with correct frequencies in ${\mathcal{O}}(10^{4})$ steps. For experiments on real image datasets, we consider a mixture of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. flipping the images (white background and black digits/objects) with probability 0.5. Following from [4], we trained a Noise Conditional Score Network (NCSN) to estimate the score function. Our numerical results indicate that Chained-LD was capable of finding both modes regardless of initialization. We summarize the contributions of this work as follows:
  • Theoretically studying the mode-seeking properties of vanilla Langevin dynamics,
  • Proposing Chained Langevin Dynamics (Chained-LD), which decomposes the sample into patches and sequentially generates each patch conditioned on previous patches,
  • Providing a theoretical analysis of the convergence behavior of Chained-LD,
  • Numerically comparing the mode-seeking properties of vanilla and chained Langevin dynamics.
@@ 2 Related Works
Langevin Dynamics: The convergence guarantees for Langevin diffusion, a continuous version of Langevin dynamics, are classical results extensively studied in the literature [12, 13, 14, 15]. Langevin dynamics, also known as Langevin Monte Carlo, is a discretization of Langevin diffusion typically modeled as a Markov Chain Monte Carlo [16]. For unimodal distributions, e.g., the probability density function that is log-concave or satisfies log-Sobolev inequality, the convergence of Langevin dynamics is provably fast [17, 18, 19]. However, for multimodal distributions, the non-asymptotic convergence analysis is much more challenging [20]. [2] gave an upper bound on the convergence time of Langevin dynamics for arbitrary non-log-concave distributions with certain regularity assumptions, which, however, could be exponentially large without imposing more restrictive assumptions. [3] studied the special case of a mixture of Gaussians of equal variance and provided an analysis of sampling from general non-log-concave distributions.
Mode-Seekingness of Langevin Dynamics: The investigation of the mode-seekingness of generative models starts with different generative adversarial network (GAN) [21] model formulations and divergence measures, from both the practical [22, 23] and theoretical [24, 7, 8] perspectives. In the context of Langevin dynamics, mode-seekingness is closely related to a lower bound on the transition time between two modes, e.g., two local maximums. [9, 10, 11] studied the mean hitting time of the continuous Langevin diffusion. [3] proved the existence of a mixture of two Gaussian distributions whose covariance matrices differ by a constant factor, Langevin dynamics cannot find both modes in polynomial time.
Score-based Generative Modeling: A central task in unsupervised learning involves learning the underlying probability distribution of training data and efficiently generating new samples from the distribution. Since [25] proposed sliced score matching which can train deep models to learn the score functions of implicit probability distributions on high-dimensional data, score-based generative modeling (SGM) has been going through a spurt of growth. Annealed Langevin dynamics [4] estimates the noise score of the probability density perturbed by Gaussian noise and utilizes Langevin dynamics to generate samples from a sequence of decreasing noise scales. [26] conducted an analysis of the effect of noise levels on the performance of annealed Langevin dynamics. Denoising diffusion probabilistic model (DDPM) [27] incorporates a step-by-step introduction of random noise into data, followed by learning to reverse this diffusion process in order to generate desired data samples from the noise. [28] unified anneal Langevin dynamics and DDPM via a stochastic differential equation. A recent line of work focuses on the non-asymptotic convergence guarantees for SGM with an imperfect score estimation under various assumptions on the data distribution [29, 30, 31, 32, 33, 34, 35]. [36] also investigated the KL convergence guarantees for score-based diffusion models. We highlight that a key difference between SGM and our theoretical analysis is that we assume the sampler has direct access to the true score function, whereas SGM typically focuses on learning the score function from training data.
@@ 3.1 Langevin Dynamics
Generative modeling aims to produce samples such that their distribution is close to the underlying true distribution $P$. For a continuously differentiable probability density $P({\mathbf{x}})$ on ${\mathbb{R}}^{d}$, its score function is defined as the gradient of the log probability density function (PDF) $\nabla_{\mathbf{x}}\log P({\mathbf{x}})$. Langevin diffusion is a stochastic process defined by the stochastic differential equation (SDE)
@@
where ${\mathbf{w}}_{t}$ is the Wiener process on ${\mathbb{R}}^{d}$. Langevin dynamics, a discretization of the SDE for $T$ iterations, is applied to generate samples. Each iteration of Langevin dynamics is defined as
@@
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. It has been widely recognized that Langevin diffusion could take exponential time to mix without additional assumptions on the probability density [9, 10, 11, 2, 3]. To combat the slow mixing, [4] proposed annealed Langevin dynamics by perturbing the probability density with Gaussian noise of variance ${\sigma}^{2}$, i.e.,
@@
and running Langevin dynamics on the perturbed data distribution $P_{{\sigma}_{t}}({\mathbf{x}})$ with gradually decreasing noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[T]}$, i.e.,
@@
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. When the noise level ${\sigma}$ is vanishingly small, the perturbed distribution is close to the true distribution, i.e., $P_{\sigma}({\mathbf{x}})\approx P({\mathbf{x}})$.
@@
In our theoretical analysis, we assume the sampler has access to the true score function $\nabla_{\mathbf{x}}\log P_{{\sigma}}({\mathbf{x}})$. In some realistic scenarios such as image datasets, since we do not have direct access to the (perturbed) score function, [4] proposed the Noise Conditional Score Network (NCSN) ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})$ to jointly estimate the scores of all perturbed data distributions, i.e., $\forall{\sigma}\in\left\{{{\sigma}_{t}}\right\}_{t\in[T]},\;{\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})\approx\nabla_{\mathbf{x}}\log P_{{\sigma}}({\mathbf{x}})$.

### v3

# On the Hardness of Sampling from Mixture Distributions via Langevin Dynamics

@@ Abstract
The Langevin Dynamics (LD), which aims to sample from a probability distribution using its score function, has been widely used for analyzing and developing score-based generative modeling algorithms. While the convergence behavior of LD in sampling from a uni-modal distribution has been extensively studied in the literature, the analysis of LD under a mixture distribution with distinct modes remains underexplored in the literature. In this work, we analyze LD in sampling from a mixture distribution and theoretically study its convergence properties. Our theoretical results indicate that for general mixture distributions of sub-Gaussian components, LD could fail in finding all the components within a sub-exponential number of steps in the data dimension. Following our result on the complexity of LD in sampling from high-dimensional variables, we propose Chained Langevin Dynamics (Chained-LD), which divides the data vector into patches of smaller sizes and generates every patch sequentially conditioned on the previous patches. Our theoretical analysis of Chained-LD indicates its faster convergence speed to the components of a mixture distribution. We present the results of several numerical experiments on synthetic and real image datasets, validating our theoretical results on the iteration complexities of sample generation from mixture distributions using the vanilla and chained LD algorithms.
@@ 1 Introduction
Langevin dynamics (LD) is a well-established methodology with a wide range of applications to various areas, including Bayesian learning (welling2011bayesian, ), non-convex optimization (raginsky2017non, ; xu2018global, ), and molecular-dynamics simulations (paquet2015molecular, ; gottwald2015parametrizing, ). The LD sampling approach leverages the score function of a probability density function (PDF) $P({\mathbf{x}})$, defined as the gradient of the PDF logarithm $\nabla\log P({\mathbf{x}})$, to perform the following iterative process whose output follows the probability model characterized by $P({\mathbf{x}})$
${\mathbf{x}}_{t}={\mathbf{x}}_{t-1}+\frac{\delta_{t}}{2}\nabla_{\mathbf{x}}\log P({\mathbf{x}}_{t-1})+\sqrt{\delta_{t}}\bm{\epsilon}_{t},$
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. Recently, the LD sampling methodology has found central applications in generative modeling tasks, such as image generation (song2019generative, ; song2020score, ), adversarial training (kamalaruban2020robust, ; srinivasan2021robustifying, ), and imitation learning (reuss2023goal, ; pearce2023imitating, ), which have inspired many theoretical and empirical studies of the LD methodology.
Specifically, several references (durmus2017nonasymptotic, ; dalalyan2017theoretical, ; cheng2018underdamped, ; cheng2018convergence, ) have studied the convergence properties of the LD sampling process to characterize the iteration complexity of LD sampling from the target PDF $P(\mathbf{x})$. The existing theoretical results mostly focus on demonstrating the satisfactory and speedy convergence of LD assuming a unimodal target distribution consisting of only one distribution component. However, the work of Song and Ermon (song2019generative, ) has highlighted examples of mixture distributions with multiple separated modes where the vanilla LD sampling struggles in capturing the mode frequencies correctly. This reference (song2019generative, ) proposes a variant of LD, called annealed Langevin dynamics (Annealed-LD), to address the challenges of LD in sampling from a mixture distribution.
While the Annealed-LD sampling approach provides a satisfactory solution to cover the modes of a mixture model, the theoretical analysis of LD sampling applied to mixture distributions has remained underexplored in the literature. However, a reliable application of LD methods to sample from real-world distributions requires a more solid understanding of their convergence behavior for a target multi-modal distribution, which is commonly present in real-world data due to the different background features of real-world objects and phenomena.
In our work, we aim to study the convergence properties of the LD framework in sampling from a mixture distribution. As displayed in Figure 1, we observe that the convergence of LD to capture all three underlying modes becomes more challenging when the dimension $d$ of the sampling space is growing. Our main theoretical result provides a family of mixture distributions, where the LD framework is unlikely to find all the mixture components within a sub-exponential number of iterations in the data dimension $d$.
Specifically, we consider mixture distributions with a low-probability yet high-variance in-between mode, which we refer to as the zeroth mode $P^{(0)}$ (illustrated in Figure 2). Despite a significantly smaller probability mass compared to the other low-variance modes, the in-between mode $P^{(0)}$ surrounds the other low-variance modes and fills the space between them. As a result, Mode 0 dominates the score function in the low-density region, disrupting and slowing down the convergence of the noisy local search in LD to the low-variance modes with greater probability masses.
(a) Mixture target distribution
(b) LD samples (scalar), $d=1$
(c) LD samples along $\overrightarrow{\bm{1}_{d}}$, $d=10$
Figure 1: Samples by Langevin dynamics from a mixture target distribution $P=0.1{\mathcal{N}}(\bm{0}_{d},10{\bm{I}}_{d})+0.45{\mathcal{N}}(5\cdot\bm{1}_{d},{\bm{I}}_{d})+0.45{\mathcal{N}}(-5\cdot\bm{1}_{d},{\bm{I}}_{d})$ with data dimensions $d=1$ and $d=10$. The samples are initialized using ${\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$, and Langevin dynamics is applied for $T=10^{4}$ iterations. The histogram is plotted by sampling $10^{4}$ vectors ${\mathbf{x}}$ and projecting them along the mean vector $\bm{1}_{d}$. Stepsizes $\delta_{t}$ are selected following (song2019generative, ).
To mitigate the exponential iteration complexity, we introduce a complementary method, Chained Langevin Dynamics (Chained-LD), with convergence guarantees in a polynomial number of iterations. Following our theoretical results on the role of high dimensionality in the convergence of LD, we propose applying dimensionality reduction through the Chain Rule: for ${\mathbf{x}}=[x^{1},x^{2},\cdots,x^{d}]\in{\mathbb{R}}^{d}$,
$P({\mathbf{x}})=P(x^{1})P(x^{2}|x^{1})\cdots P(x^{d}|x^{1},\cdots,x^{d-1}).$
Chained-LD sequentially samples every element ${\mathbf{x}}^{i}$ for all $i\in[d]$ from the conditional distribution given previous elements, i.e., $P(x^{i}\mid x^{1},\cdots x^{i-1})$. Therefore, Chained-LD reduces the effective dimensionality of the sampled variable, which can accelerate the search for missing modes in sampling from a mixture distribution. Furthermore, for mixture distributions $P$ such that $-\log P(x^{i}|x^{1},\cdots,x^{i-1})$ is $L_{Q}$-smooth and $m_{Q}$-strongly convex for $|x^{i}|>R_{Q}$, we theoretically show that Chained-LD converges to the target distribution within $\varepsilon$ total variation distance in ${\mathcal{O}}\left(\frac{L_{Q}^{2}d^{3}}{m_{Q}^{2}\varepsilon^{2}}\exp(32L_{Q}R_{Q}^{2})\log\frac{d^{3}}{\varepsilon^{2}}\right)={\mathcal{O}}\left(\frac{d^{3}}{\varepsilon^{2}}\log\frac{d^{3}}{\varepsilon^{2}}\right)$ iterations.
@@
Finally, we present the results of several numerical experiments to validate our theoretical findings. In synthetic experiments, we consider high-dimensional Gaussian mixture models, where LD could not find all components within a million steps, whereas Chained-LD could capture all components with correct frequencies in ${\mathcal{O}}(10^{4})$ steps. Also, we test the application of Chained-LD as a sampling algorithm in score-based generative modeling for an underlying mixture distribution. In the case of a mixture of original images from the MNIST/Fashion-MNIST dataset (black background and white digits/objects) and flipped images (white background and black digits/objects), our numerical results suggest that Chained-ALD could find both the modes in ${\mathcal{O}}(10^{5})$ iterations. We summarize the contributions of this work as follows:
  • Analyzing the iteration complexity of Langevin dynamics under high-dimensional mixture distributions,
  • Proposing Chained Langevin Dynamics (Chained-LD) with sequential sampling to improve LD’s convergence in sampling from mixture distributions,
  • Providing a theoretical analysis of the convergence of Chained-LD,
  • Presenting numerical results validating our theoretical findings on the convergence of LD and Chained-LD.
@@ 2 Related Works
Convergence Guarantees of Langevin Dynamics: The convergence guarantees of Langevin diffusion, a continuous version of Langevin dynamics, are classical results that have been extensively studied in the literature (bhattacharya1978criteria, ; roberts1996exponential, ; bakry1983diffusions, ; bakry2008simple, ). Langevin dynamics, also known as Langevin Monte Carlo, is a discretization of Langevin diffusion typically modeled as a Markov Chain Monte Carlo (MCMC) method. For uni-modal distributions, e.g., log-concave probability density functions, the convergence of Langevin dynamics is provably fast (dalalyan2017theoretical, ; durmus2017nonasymptotic, ; cheng2018underdamped, ; cheng2018convergence, ). However, for multi-modal distributions, the non-asymptotic convergence analysis becomes significantly more challenging. (lee2018beyond, ) studied Langevin dynamics under mixtures of Gaussians with equal variance and showed that the iteration complexity of Langevin dynamics is $\text{poly}(d,1/\varepsilon)$. For more general distributions, (cheng2018sharp, ) and (ma2019sampling, ) analyzed target distributions $p$ that are strongly log-concave outside of a region of radius $R$, proving that the iteration complexity of Langevin dynamics is $\exp(cR^{2})\text{poly}(d,1/\varepsilon)$, which can become exponential in $d$ when the radius $R$ scales as ${\mathcal{O}}(\sqrt{d})$.
Hardness of Langevin Dynamics in Mixture Distributions: For continuous Langevin diffusion, (bovier2002metastability, ; bovier2004metastability, ; gayrard2005metastability, ) studied the mean hitting time and provided a lower bound on the transition time between two modes, e.g., two local maxima. In the context of Langevin dynamics, (lee2018beyond, ) proved the existence of a mixture of two Gaussian distributions with covariance matrices differing by a constant factor, wherein Langevin dynamics cannot find both modes in polynomial time. (song2019generative, ) studied the slow mixing and incorrect relative weight recovery of Langevin dynamics in bi-modal distributions separated by low-density regions. Additionally, (song2020improved, ) studied the role of noise levels in annealed Langevin dynamics, showing their effect on sample diversity in multi-modal distributions.
Connections between Langevin Dynamics and Score-based Generative Modeling: Langevin dynamics and its annealed variant serve as the backbone of score-based generative modeling, which aims to learn the underlying probability distribution of training data and efficiently generate new data from the learned distribution. (song2019generative, ) proposed learning Noise Conditional Score Networks (NCSN) via score matching to estimate the perturbed score function of the underlying distribution from training data and applied annealed Langevin dynamics with NCSN as the sampling method. (song2020score, ) unified anneal Langevin dynamics and Denoising diffusion probabilistic modeling (DDPM) (ho2020denoising, ) via a stochastic differential equation (SDE) and proposed utilizing score-based Markov Chain Monte Carlo (MCMC) approaches, e.g., Langevin dynamics, to sample from the SDE.
@@ 3 Preliminaries
(a) Symmetric 3-Gaussian model
(b) Multimodal distribution
Figure 2: Our analyzed mixture distribution possessing the in-between mode $P^{(0)}$. $P^{(0)}$ is supposed to contain a minor probability mass, yet with a significantly higher variance than the other modes $P^{(1)},\ldots,P^{(k)}$.
@@ 3.1 Langevin Dynamics
Langevin dynamics aims to produce samples such that their distribution is close to the underlying true distribution $P$. For a continuously differentiable probability density $P({\mathbf{x}})$ on ${\mathbb{R}}^{d}$, its score function is defined as the gradient of the log probability density function (PDF) $\nabla_{\mathbf{x}}\log P({\mathbf{x}})$. Langevin diffusion is a stochastic process defined by the stochastic differential equation (SDE)
@@
where ${\mathbf{w}}_{t}$ is the Wiener process on ${\mathbb{R}}^{d}$. Langevin dynamics, a discretization of the SDE for $T$ iterations, is applied to sample from the target distribution. Each iteration of Langevin dynamics is defined as
@@
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. It has been widely recognized that the continuous Langevin diffusion could take an exponential time to mix without additional assumptions on the probability density (bovier2002metastability, ; bovier2004metastability, ; gayrard2005metastability, ). To combat the slow mixing, (song2019generative, ) proposed annealed Langevin dynamics by perturbing the probability density with Gaussian noise of variance ${\sigma}^{2}$, i.e.,
@@
and applying Langevin dynamics on the perturbed data distribution $P_{{\sigma}_{t}}({\mathbf{x}})$ with gradually decreasing noise levels ${\sigma}_{1}\geq{\sigma}_{2}\geq\cdots\geq{\sigma}_{T}$, i.e.,
@@
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. When the noise level ${\sigma}$ is vanishingly small, the perturbed distribution is close to the true distribution, i.e., $\lim_{{\sigma}\to 0}P_{\sigma}({\mathbf{x}})\approx P({\mathbf{x}})$.
@@
In our theoretical analysis, we assume the sampler has access to the underlying score function $\nabla_{\mathbf{x}}\log P_{{\sigma}}({\mathbf{x}})$. For generative modeling tasks in real-world datasets, since we do not have direct access to the (perturbed) score function, (song2019generative, ) proposed the Noise Conditional Score Network (NCSN) ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})$ to jointly estimate the scores of all perturbed data distributions, i.e., $\forall{\sigma}\in\left\{{{\sigma}_{t}}\right\}_{t\in[T]},\;{\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})\approx\nabla_{\mathbf{x}}\log P_{{\sigma}}({\mathbf{x}})$.